If a person rolls doubles when tossing two dice, the roller profits $50, that is, the roller receives $50 more than the cost of the game. If the game is fair, how much should the person pay to play the game?
Answer #1

probability of getting doubles is 6/36 = 1/6

As the game is fair, the total value should be zero

1/6*50 - 5/6*x = 0
x = 50/6 * 6/5
x = 10

Hence, $10 should be paid